# Semantic Residual Prompts for Continual Learning

### v1

@@ 3.4 Discussion and comparisons to related works
The dominant contribution of our work is the use of a foundation model to increase the stability of the prompt selection. Specifically, we regard a prompting strategy as stable if, for a given task, the query-key selection strategy retrieves the set of prompts relevant to that task, and not those introduced for subsequent tasks. In the following, we are dissecting why existing approaches overlook stability, leading to prompts of one task being wrongly used for other tasks, and how we exploit the CLIP model to further it.
On the stability of prompting-based CL methods. Like our method, existing approaches such as L2P, DualPrompt, and CODA-Prompt employ a query-key matching mechanism to fetch the prompts utilized for adapting the pre-trained ViT. Notably, they learn new keys from scratch when a new task is presented, and define a customized loss term to incentives the learned keys to be aligned with the visual queries of the current task (extracted using the pre-trained ViT itself). However, these learned keys often lack a clear semantic interpretation, and could be theoretically used to represent distant concepts.
@@
In this respect, we argue that learning keys as such is susceptible to forgetting, and to an interference between prior and subsequent tasks. Indeed, since keys are continuously adjusted to align with the queries of the current task, they may drift and no longer match with the queries of earlier tasks. Or, when introducing novel keys, they could interfere with those already present in the pool, and be selected to classify examples of previous tasks.
@@
We assess such an intuition by analyzing the prompts retrieved by different methods – i.e., DualPrompt, CODA-Prompt, and our approach STAR-Prompt– on the Split Imagenet-R dataset. In detail, Fig. 3, shows the proportion of testing samples of the first task $D_{1}$ retrieving the right prompts — namely those corresponding to prompts learned during the first task $D_{1}$. As can be seen, the precision of the selection strategies of both DualPrompt and CODA-Prompt decreases as the tasks progress, indicating that more and more examples of $D_{1}$ use prompts of subsequent tasks. Our approach is instead much more stable, as also evident in Fig. 3, which depicts the full-task confusion matrices computed after the final task. While the confusion matrix of STAR-Prompt is almost diagonal, the prompt selection of DualPrompt and CODA-Prompt yields significant task confusion. As outlined in the suppl. materials, we verified these outcomes also on other domains (i.e., in aerial and medical datasets).
@@
The reason at the base of the stability of our STAR-Prompt is twofold. On the one hand, we do not start from scratch when learning the keys, but we fine-tune the CLIP text encoder, which allows to devote distinct learnable parameters to distinct classes. Such an approach promotes a greater modularity, and has been already acknowledged as more stable by the authors of PromptFusion [5]. Differently from them, however, we exploit the stability of the text encoder not for direct prediction, but to define a second pool codebook of prompts.
@@
On the other hand, we provide explicit supervision (Eq. 8 and Eq. 15) to enforce class-driven separation in the key space. This differs from other CL prompting approaches, which settle to weaker constraints (e.g., relying on orthogonality or aligning the query with the closest key in the pool). Hence, since first-level $\boldsymbol{w}_{c}$ keys are both stable and separated, the second prompts $Q_{c}$ can learn class-specific information, reducing the interference among classes/tasks.
@@
Prompting mechanism. In VPT, PromptFusion, and most prompt-tuning methods, the prompt tokens are concatenated with the input sequence before each MSA layer. Another common solution is prefix tuning (Sec. 2), which intervenes only on the keys and the values of the MSA layer. Since the frozen MSA layer weighs the importance of each prompt token, we argue that there is no barrier preventing some of these tokens from being disregarded.
Conversely, our semantic residuals are added after the MSA layer (see Eq. 7). As we sum the prompts directly to the input of the MLP layers, the corresponding hint from the CLIP encoders cannot be ignored. This is similar to the shifting parameters of FiLM layers [30] used to adapt the Batch Norm [16] of CNNs. However, we do not have a parameter “generator” network [30], and the residuals for the Transformer MLPs are computed using CLIP-guided shifting vectors.
@@ 3.5 Multi-Modal Generative Replay
1:Index of the current task $t\in{1,\dots,T}$, the orthogonality loss weighting coefficient $\lambda$, the number of $M$ Gaussian components, the number of $E_{1}$ and $E_{2}$ training epochs with real and generated samples (respectively), learning rate $lr$.
2:First stage $\rightarrow$ goal: learning first-level prompts $\boldsymbol{p}_{c}|_{c\in\{1,...,\lvert{\cal Y}_{t}\lvert\}}$ for each class $c\in{\cal Y}_{t}$ of the current task $t$.
@@
6:   $\boldsymbol{p}_{c}\xleftarrow{}\boldsymbol{p}_{c}-lr\cdot\nabla_{\boldsymbol{p}_{c}|_{c\in{\cal Y}_{t}}}{\cal L}_{\operatorname{CE}}+\lambda{\cal L}_{\operatorname{OP}}$
@@
13:   $\boldsymbol{p}_{c}\xleftarrow{}\boldsymbol{p}_{c}-lr\cdot\nabla_{\boldsymbol{p}_{c}|_{c\in{\cal Y}_{t}}}{\cal L}_{\operatorname{GR}}^{P}$
@@
19:   $\mathcal{L}_{\operatorname{OQ}}^{P}\xleftarrow{}$ use Eq. 12
@@
23:$\mathcal{H}_{c}\xleftarrow{}\operatorname{EM}(\{f(x_{i};R)\}_{x,y=c\in D_{t}})\ \forall c\in{\cal Y}_{t}$
@@
Since for each task we train only the prompts of the current task, to further reduce possible inter-task inferences, we adopt a generative replay strategy inspired by SLCA [48] (Sec. 2). In [48], class-specific features are used to compute the mean and the covariance matrix of a Gaussian representing the class distribution. Then, subsequent tasks use this Gaussian to generate $n=256$ synthetic features of that class, as a form of generative replay. We extend this idea by introducing, in both stages of our training, a Mixture of Gaussians (MoGs) representation of the feature distribution. Specifically, in the first stage, for each class $c$, we fit a MoGs $\mathcal{G}_{c}$ on the CLIP visual features $\boldsymbol{z}_{i}=E_{vis}(x_{i})\ \forall\ (x_{i},y_{i}=c)\in D_{t}$:
@@
The parameter vector $\boldsymbol{\phi_{c}}$ contains the $M$ means, covariances and corresponding Gaussian component weights, and it is estimated by Expectation-Maximization (EM). $\mathcal{G}_{c}$ is then used in the subsequent tasks $t^{\prime}$ ($t^{\prime}\geq t$) to generate $n=256$ synthetic features $\tilde{\boldsymbol{z}}_{j}$ of class $c$. Using this generative replay technique, we obtain:
@@
Similarly, in the second training stage, we use the ViT features $\boldsymbol{e}_{L}^{\operatorname{CLS}}=f(x_{i};R)$ of the current task to estimate the parameters of a second MoGs, indicated with $\mathcal{H}_{c}$, as for Eq. 13. In the rehearsal phase, we replace $\boldsymbol{e}_{L}^{\operatorname{CLS}}$ in Eq. 10 with $n$ synthetic features sampled from each $\mathcal{H}_{c}$, and, analogously to Eqs. 15 and 14, we compute ${\cal L}_{\operatorname{GR}}^{Q}$ (details in the suppl. materials).
The whole training algorithm is shown in Algorithm 1. For ease of presentation, we do not split the dataset in mini-batches when computing a Gradient Descent (GD) step. The algorithm shows that, in the first training stage, only the first-level prompts are updated ($\boldsymbol{p}_{c},c\in{\cal Y}_{t}$, see lines 5 and 11). Conversely, in the second stage, we train the second-level prompts ($Q_{c}$), the query weighting parameters ($A_{c}$), and the linear layer of $g(\cdot)$. Specifically, in the second-stage rehearsal phase, we update the heads of all the tasks observed so far (i.e., $\theta_{t^{\prime}\leq t}$).
@@ 4 Experiments
Datasets.  We provide an evaluation encompassing a variety of benchmarks. We follow the current literature dealing with pre-trained CL models [46, 40, 48, 5] and cover conventional image dataset such as CIFAR-100 and Imagenet-R. Then, we assess the adaptability of these pre-trained methods by moving to settings with decreasing domain similarity w.r.t. the Image-Net pre-training [28, 8]. In particular, we test on the following domains (more details can be found in suppl. materials):
@@
  • Specialized domain: following [48], we adopt Split Cars-196 [21] and Split CUB-200 [42] as fine-grained classification scenarios. Classes are split across $10$ tasks.
@@
  • Medical domain: we adopt three settings, ranging from plant to human diseases. Split CropDiseases [15] regards healthy/infected leaves with $7$ tasks of $5$ classes each. In Split ISIC [7], images depicting $6$ skin diseases are split into $3$ tasks. Split ChestX [45] consists of chest X-ray images ($2$ tasks with $3$ classes each).
Metrics.  We report the average accuracy of each task, computed after the end of the last one (Final Average Accuracy – FAA). We also report the Final Forgetting [4] in suppl. materials. Each experiment is repeated three times with varying class orders [48]. We hence provide both the mean and the standard deviation of the FAA.
Implementation details. We train our model with Adam [19] with a learning rate of $0.001$. The number of epochs is $50$, with some exceptions due to the varying dataset sizes (see suppl. materials). The size of each training mini-batch is fixed at $16$ for Split Imagenet-R and $128$ for the other datasets. For a fair comparison, we maintain the same number of epochs for all the competitors but search their optimal hyper-parameters (including the learning rate and the optimizer) to guarantee their best performance. Due to space limits, we report all method-specific hyper-parameters in suppl. materials.
@@
Training-free baselines and lower/upper bounds.  We integrate our analysis with the zero-shot testing of CLIP [32] (Zero-shot CLIP), which uses hand-crafted textual prompts (e.g., “A photo of a $\operatorname{[CL-NAME]}$”) to classify a testing image through Eq. 9. Notably, the comparison between STAR-Prompt and the Zero-shot CLIP test provides an understanding of the algorithmic merits of our approach, and allows us to disentangle them from the strengths of CLIP. Futhermore, we consider a similar baseline built on top of the ImageNet pre-trained backbone. It freezes the network and simply memorize the latent features of each training example, employing a k-NN approach.
@@
Joint (STAR-Prompt) $89.49$$\pm$$0.12$ $92.37$$\pm$$0.09$ $88.65$$\pm$$0.19$ $86.37$$\pm$$0.15$ $89.22$
@@
Zero-shot CLIP [32] $82.16$ $72.62$ $64.98$ $50.53$ $67.57$
@@
SLCA (3) † [48] $77.00$$\pm$$0.33$ $\mathbf{91.53}$$\pm$$0.28$ $67.73$$\pm$$0.85$ $84.71$$\pm$$0.40$ $80.24$
STAR-Prompt $\mathbf{89.16}$$\pm$$0.38$ $90.16$$\pm$$0.34$ $\mathbf{86.50}$$\pm$$0.52$ $\mathbf{85.24}$$\pm$$0.20$ $\mathbf{87.77}$
For a thorough comparison, we also include the results of STAR-Prompt when jointly trained on all tasks in a stationary scenario (denoted as Joint (STAR-Prompt)), which serves as a performance upper bound. Moreover, as commonly done in the literature, we include the results of Joint  (ViT-B/16) and Fine-tune  (ViT-B/16), where the ViT-B/16 architecture is fine-tuned on all tasks respectively jointly and incrementally with no countermeasure to forgetting.
@@ 4.1 Comparison with the State of the Art
Joint (STAR-Prompt) $97.19$ $96.50$ $99.28$ $79.61$ $46.07$ $83.73$
@@
Zero shot CLIP $57.74$ $59.60$ $36.77$ $34.29$ $16.81$ $41.04$
@@
STAR-Prompt $\mathbf{94.24}$ $\mathbf{92.01}$ $\mathbf{94.73}$ $\mathbf{67.39}$ $\mathbf{40.47}$ $\mathbf{77.77}$
Tab. 1 and 2 provide a detailed performance comparison. We refer the reader to the suppl. materials for a wider analysis covering the number of parameters. As can be seen, STAR-Prompt outperforms all the other approaches, usually by a large margin.
The comparison is particularly significant when referring to other CLIP-based methods (i.e., AttriCLIP, and PromptFusion), including Zero-shot CLIP. For instance, in the two fine-grained scenarios – Split Cars-196 and Split CUB-200– we get a $+21.52$ and a $+34.71$ margin over Zero-shot CLIP, respectively. We ascribe this result to the fact that fine-grained class differences (distinguishing an “Audi S4 Sedan 2007” from an “Audi S6 Sedan 2011”) are not easily captured by the pre-trained CLIP. Therefore, a mechanism devoted to plasticity as our second-level prompting stage is fundamental to address these scenarios. A similar conclusion could be achieved for both the aerial and the medical domains, where the average gap w.r.t. the Zero-shot CLIP is $+36.73$.
Remarkably, we get a $+5.84$ average boost over the runner-up method (SLCA). Such a result is even more impressive considering that, compared to the latter, STAR-Prompt includes only a fraction of its learnable parameters, making it far easier to train (see suppl. materials for more details). Only Split CIFAR-100 shows slight inferior performance, a result we ascribe to the up-sampling operation that is necessary on its images (a factor of ×$\times 7$), which introduces significant visual artifacts.
We also emphasize that STAR-Prompt closely matches the performance of its Joint upper bound, with an average margin of just $-3.96$. We feel this result as particularly promising: our solution yields an almost negligible drop in the incremental training with respect to a stationary scenario. Such an outcome holds true for a scenario demanding high adaptability as the aerial one, while the gap is a bit larger for the medical one.
We also observe that existing prompting approaches are poorly effective in aerial and medical settings. They are surpassed by replay-based approaches, which can deeper fine-tune the backbone to account for the domain shifts. Notably, STAR-Prompt stands out, suggesting it represents the best compromise between stability and plasticity.
Finally, we highlight that the overall complexity of STAR-Prompt is comparable to that of other prompt-based methods. In fact, as the class prototypes $\boldsymbol{w}_{c}$ do not depend on the input image, they can be computed once and then cached after each task.

### v3

@@ 3.4 Discussion and comparisons to related works
Our main contribution relies on the use of a foundation model to enhance the stability of the prompt selection. Indeed, we regard a prompting strategy as stable if, given a task, the query-key selection strategy retrieves the set of prompts relevant to that task, and not those introduced for subsequent tasks. In the following, we dissect why existing approaches overlook stability and how we exploit CLIP to enhance it.
On the stability of prompting-based CL methods. Like our method, existing approaches such as L2P, DualPrompt, and CODA-Prompt employ a query-key matching mechanism to fetch the prompts utilized for adapting the pre-trained ViT. Notably, they learn new keys from scratch when a new task is presented, and define a customized loss term to incentivize the learned keys to be aligned with the visual queries of the current task (extracted using the pre-trained ViT itself). However, these learned keys often lack a clear semantic interpretation, and could be theoretically used to represent distant concepts.
@@
Concerning this, we argue that learning keys as such is susceptible to interference between tasks. Indeed, since keys are continuously adjusted to align with the queries of the current task, they may drift and no longer match with the queries of earlier tasks. Or, when introducing novel keys, they could interfere with those already present in the pool, and be selected to classify examples of previous tasks. We assess such an intuition by analyzing the prompts retrieved by different methods – i.e., DualPrompt, CODA-Prompt, and our approach STAR-Prompt– on the Split Imagenet-R dataset. In detail, Fig. 3, shows the proportion of testing samples of the first task $D_{1}$ retrieving the right prompts – namely, those corresponding to prompts learned during the first task $D_{1}$. As can be seen, the precision of the selection strategies of both DualPrompt and CODA-Prompt decreases as the tasks progress, indicating that more and more examples of $D_{1}$ use prompts of subsequent tasks. Our approach is instead much more stable, as also evident in Fig. 3, which depicts the full confusion matrices computed after the final task. While the confusion matrix of STAR-Prompt is almost diagonal, the prompt selection of DualPrompt and CODA-Prompt yields significant task confusion. As outlined in the suppl. materials, we verified these outcomes also on other domains (i.e., in aerial and medical datasets).
@@
The reason at the base of the stability of our STAR-Prompt is twofold. On the one hand, we do not start from scratch when learning the keys, but we adapt the CLIP text encoder, which allows us to devote distinct learnable parameters to distinct classes. Such an approach promotes a greater modularity, and has been already acknowledged as more stable by the authors of PromptFusion [6]. Differently from them, however, we exploit the stability of the text encoder not for direct prediction, but to fetch a second codebook of prompts.
@@
On the other hand, we provide explicit supervision (Eq. 8) to enforce class-driven separation in the key space. This differs from other CL prompting approaches, which settle to weaker constraints (e.g., relying on orthogonality or aligning the query with the closest key in the pool). Hence, since $\boldsymbol{w}_{c}$ keys are both stable and separated, the second-level prompts $Q_{c}$ can learn class-specific information, reducing the interference among classes/tasks.
@@
Prompting mechanism. In most methods, prompts are prepended with the input sequence before each MSA layer. Another common solution is prefix tuning (Sec. 2), which intervenes only on the keys and the values of the MSA layer. Since the frozen MSA layer weighs the importance of each prompt, we argue that there is no barrier preventing some of these tokens from being disregarded. Conversely, our semantic residuals are added after the MSA layer (see Eq. 7). This is similar to the shifting parameters of FiLM layers [31] used to adapt the Batch Norm [17] of CNNs. However, we do not have a parameter “generator” network [31], and the residuals for the Vision Transformer MLPs are computed using CLIP-guided shifting vectors.
Computational cost. As detailed in suppl. materials, the computational demand associated with STAR-Prompt is comparable with that of most prompt-based approaches (e.g., L2P, CODA-Prompt, etc.), requiring two forward passes. The key difference is that these methods rely on the same backbone, whereas we employ two backbones with distinct parameters.
@@ 3.5 Multi-Modal Generative Replay
1:Index of the current task $D_{t},\ t\in{1,\dots,T}$, the orthogonality loss weighting coefficient $\lambda$, the number of $M$ Gaussian components, the number of $E_{1}$ and $E_{2}$ training epochs with real and generated samples (respectively), learning rate $lr$.
2:First stage $\rightarrow$ goal: learning first-level prompts $p_{c}|_{c\in\{1,...,N\}}$ for each class $c\in{\cal Y}_{t}$ of the current task $t$.
@@
6:   $p_{c}\xleftarrow{}p_{c}-lr\cdot\nabla_{p_{c}|_{c\in{\cal Y}_{t}}}{\cal L}_{\operatorname{CE}}+\lambda{\cal L}_{\operatorname{OP}}$
@@
13:   $p_{c}\xleftarrow{}p_{c}-lr\cdot\nabla_{p_{c}|_{c\in{\cal Y}_{t}}}{\cal L}_{\operatorname{GR}}^{P}$
@@
19:   $\mathcal{L}_{\operatorname{OQ}}\xleftarrow{}$ use Eq. 12
@@
23:$\mathcal{H}_{c}\xleftarrow{}\operatorname{EM}(\{f(x_{i};\mathbf{R})\}_{x,y=c\in D_{t}})\ \forall c\in{\cal Y}_{t}$
@@
In SLCA, class-specific features are used to estimate the statistics of a Gaussian representing the class distribution. Then, subsequent tasks use this Gaussian to generate synthetic features and perform generative replay. To capture multiple peaks in the distribution of each class, we extend this idea by introducing, in both stages of our training, a Mixture of Gaussians (MoGs) representation of the feature distribution. Specifically, in the first stage, for each class $c$, we fit a MoGs $\mathcal{G}_{c}$ on the CLIP visual features $\boldsymbol{z}_{i}=E_{vis}(x_{i})\ \forall\ (x_{i},y_{i}=c)\in D_{t}$:
@@
The parameter vector $\boldsymbol{\phi_{c}}$ contains the $M$ means, covariances and corresponding Gaussian component weights, and it is estimated by Expectation-Maximization (EM). $\mathcal{G}_{c}$ is then used in the subsequent tasks $t^{\prime}$ ($t^{\prime}\geq t$) to generate $n=256$ synthetic features $\tilde{\boldsymbol{z}}$ of class $c$. Using this generative replay technique, we obtain:
@@
Similarly, in the second training stage, we use the ViT features $\boldsymbol{e}_{L}^{\operatorname{CLS}}=f(x_{i};\mathbf{R})$ of the current task to estimate the parameters of a second MoGs, indicated with $\mathcal{H}_{c}$, as for Eq. 13. In the rehearsal phase, we replace $\boldsymbol{e}_{L}^{\operatorname{CLS}}$ in Eq. 10 with $n$ synthetic features sampled from each $\mathcal{H}_{c}$, and, analogously to Eqs. 15 and 14, we compute ${\cal L}_{\operatorname{GR}}^{Q}$ (details in the suppl. materials).
The whole training algorithm is shown in Algorithm 1. For ease of presentation, we do not split the dataset in mini-batches when computing a step of gradient descent. The algorithm shows that, in the first training stage, only the first-level prompts are updated (i.e., $p_{c},c\in{\cal Y}_{t}$). Conversely, in the second stage, we train the second-level prompts $Q_{c}$, the query weighting parameters $A_{c}$, and the linear layer of $g_{{\theta}_{t}}(\cdot)$. Specifically, in the second-stage rehearsal phase, we update the heads of all the tasks observed so far (i.e., $\theta_{t^{\prime}\leq t}$).
@@ 4 Experiments
We evaluate our approach on varying benchmarks. We follow the current literature dealing with pre-trained CL models [48, 41, 50, 6] and cover conventional image datasets such as CIFAR-100 and Imagenet-R. Then, we assess the adaptability of these pre-trained methods by moving to settings with decreasing domain similarity w.r.t. the Image-Net pre-training [29, 9]. In particular, we test on the following domains (more details can be found in suppl. materials):
@@
  • Specialized domain: following [50], we adopt Split Cars-196 [22] and Split CUB-200 [44] as fine-grained scenarios. Classes are split across $10$ tasks.
@@
  • Medical domain: we adopt three settings, ranging from plant to human diseases. Split CropDiseases [16] regards healthy/infected leaves with $7$ tasks/$5$ classes each. In Split ISIC [8], images of $6$ skin diseases are split into $3$ tasks. Split ChestX [47] consists of chest X-ray images ($2$ tasks/$3$ classes each).
Metrics.  We report the Final Average Accuracy (FAA) – computed after the end of the last task – and the Final Forgetting [5] in suppl. materials. Each experiment is repeated three times with varying class orders [50]. We hence provide both the mean and the standard deviation of the FAA.
Implementation details. We train our model with Adam [20] with a learning rate of $0.001$. The number of epochs is set according to the varying dataset sizes (see suppl. materials), while we adopt a batch size of $16$ for Split Imagenet-R and $128$ for the other datasets. For a fair comparison, we maintain the same number of epochs for all the competitors and search their optimal hyper-parameters (including learning rate and optimizer) to guarantee the best performance. These configurations are reported in suppl. materials. In line with most CL literature, we perform inference in an instance-wise setup, where each prompt is selected independently for each sample of the batch (more in suppl. materials).
@@
Training-free baselines and lower/upper bounds.  We integrate our analysis with the zero-shot testing of CLIP [33] (Zero-shot CLIP), which uses hand-crafted textual prompts (e.g., “A photo of a $\operatorname{[CLS-NAME]}$”) to classify a testing image through Eq. 9. Notably, the comparison between STAR-Prompt and the Zero-shot CLIP test provides an understanding of the algorithmic merits of our approach, and allows us to disentangle them from the strengths of CLIP. Futhermore, we consider a similar baseline built on top of the ImageNet pre-trained backbone. It freezes the network and simply memorize the latent features of each training example, employing a k-NN approach.
@@
Joint (STAR-Prompt) $90.03$$\pm$$0.33$ $92.32$$\pm$$0.28$ $89.00$$\pm$$0.33$ $85.64$$\pm$$0.30$ $89.25$
@@
Zero-shot CLIP [33] $85.36$ $73.27$ $76.46$ $61.14$ $74.06$
@@
SLCA (3) † [50] $77.00$$\pm$$0.33$ $\mathbf{91.53}$$\pm$$0.28$ $67.73$$\pm$$0.85$ $\mathbf{84.71}$$\pm$$0.40$ $80.24$
STAR-Prompt $\mathbf{89.83}$$\pm$$0.04$ $90.12$$\pm$$0.32$ $\mathbf{87.62}$$\pm$$0.20$ $84.10$$\pm$$0.28$ $\mathbf{87.92}$
For a thorough comparison, we include the results of STAR-Prompt when trained on all tasks in a stationary scenario – Joint (STAR-Prompt) – which serves as an upper bound. Moreover, we include Joint  (ViT-B/16) and Fine-tune  (ViT-B/16), where a ViT-B/16 is fine-tuned on all tasks respectively jointly and incrementally with no countermeasure to forgetting.
@@ 4.1 Comparison with the State of the Art
Joint (STAR-Prompt) $97.21$ $96.41$ $99.19$ $78.25$ $45.36$ $83.28$
@@
Zero shot CLIP $54.50$ $63.15$ $27.58$ $29.14$ $26.27$ $40.13$
@@
STAR-Prompt $\mathbf{93.70}$ $\mathbf{92.28}$ $\mathbf{94.92}$ $\mathbf{66.67}$ $\mathbf{41.85}$ $\mathbf{77.88}$
As shown in Tab. 1 and 2, STAR-Prompt outperforms all the other approaches on average. The comparison is particularly significant when referring to other CLIP-based methods (e.g. AttriCLIP and PromptFusion). For instance, in the two fine-grained scenarios – Cars-196 and CUB-200– we get a $+11.16$ and a $+22.96$ margin over Zero-shot CLIP, respectively. We ascribe this result to the fact that fine-grained class differences (distinguishing a “red headed woodpecker” from a “red bellied woodpecker”) are not easily captured by the pre-trained CLIP. Therefore, a mechanism devoted to plasticity as our second-level prompting stage becomes fundamental. A similar conclusion could be drawn for the aerial and medical domains, where the average gap w.r.t. the Zero-shot CLIP is $+37.75$.
Remarkably, we get a $+5.96$ average boost over the runner-up method (SLCA), which is remarkable considering that STAR-Prompt has only a fraction of its learnable parameters (see suppl. materials). We also emphasize that STAR-Prompt closely matches the performance of its Joint upper bound, with an average margin of just $-3.59$. This result looks particularly promising: our solution yields an almost negligible drop in performance w.r.t. a stationary scenario. Such an outcome holds true for both aerial and medical scenarios, which demand high adaptability.
Moreover, existing prompting approaches struggle in aerial and medical settings. They are surpassed by replay-based approaches, which can deeper fine-tune the backbone to account for domain shifts. Notably, STAR-Prompt stands out as the best compromise between stability and plasticity.